$SPAD/src/input drawalg.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- TopLevelDrawFunctionsForAlgebraicCurves example
--Plotting Plane Algebraic Curve
draw(y**2 + y - (x**3 - x) = 0, x, y, range == [-2..2,-2..1])

———-

References^{[1] nothing}

4